# VADA: a Data-Driven Simulator for Nanopore Sequencing

### v1

@@ Abstract
Nanopore sequencing offers the ability for real-time analysis of long DNA sequences at a low cost, enabling new applications such as early detection of cancer. Due to the complex nature of nanopore measurements and the high cost of obtaining ground truth datasets, there is a need for nanopore simulators. Existing simulators rely on handcrafted rules and parameters and do not learn an internal representation that would allow for analysing underlying biological factors of interest. Instead, we propose VADA, a purely data-driven method for simulating nanopores based on an autoregressive latent variable model. We embed subsequences of DNA and introduce a conditional prior to address the challenge of a collapsing conditioning. We introduce an auxiliary regressor on the latent variable to encourage our model to learn an informative latent representation. We empirically demonstrate that our model achieves competitive simulation performance on experimental nanopore data. Moreover, we show we have learned an informative latent representation that is predictive of the DNA labels. We hypothesize that other biological factors of interest, beyond the DNA labels, can potentially be extracted from such a learned latent representation.
@@ 1 Introduction
To condition our model in accordance with the physical properties of nanopore sequencing, we represent the DNA sequence by embedding k-mers of DNA, as done in other machine learning tasks in this domain [16]. Initial empirical results showed that a straightforward approach to conditioning the latent variable model results in a conditioning collapse, where the model ignores the conditioning and produces bad samples unrelated to the DNA sequence. To overcome this problem, and effectively condition our model on a sequence of DNA, we introduce a conditional prior distribution on the latent space. Moreover, we introduce an auxiliary regressor on the latent space to encourage the model to learn a latent representation that contains information about the DNA sequence.
@@ 2.1.2 DNA representation
Multiple DNA bases influence each nanopore current measurement. Consequently, as mentioned earlier, we represent $y_{t}$ as a k-mer containing $k=5$ DNA bases. The embedding layer $f_{\theta}$ processes each $y_{t}$ in the window independently, producing embeddings $e^{t:t+\delta}=f_{\theta}(y^{t:t+\theta})$.
@@ 2.1.4 Informative latent space
Our model inherently learns a latent representation of the nanopore currents via the latent variable $z$ and is further encouraged to encode information about the DNA sequence into $z$ through the use of the conditional prior [7]. Nevertheless, there is no guarantee that $z$ captures the true underlying sources of variation, such as the DNA sequence that was in the pore. Therefore, to encourage our model to learn a representation that contains information about the DNA sequence, we make use of an auxiliary regressor during training, as done in other VAE’s that aim to learn an informative latent space, such as DIVA [7]. Specifically, during training, this regressor processes a latent sample $z$ to predict the aggregated embedding, $y^{\prime}$, that was used to condition the prior, as $\hat{y}^{\prime}=r_{\theta}(z)$. The auxiliary regressor is included in the visualization of the training procedure (Fig. 3).
Figure 3: VADA training overview.
@@ 2.1.5 Training
Following the VAE framework [9] we utilize an approximate posterior $q_{\theta}(z|x^{t:t+\delta})$ during training. We optimize our model using a modified $\beta$-VAE loss [6, 7], where Mean Squared Error (MSE) is used to optimize the auxiliary regressor. The contribution of the MSE to the overall loss is scaled using a hyperparameter $\beta_{\text{aux}}$. The complete loss term is given by Equation 2.
$\mathcal{L}(x^{t:t+\delta},x^{t-\delta:t},y^{t:t+\delta})={}$ $\mathbb{E}_{q_{\theta}(z|x^{t:t+\delta})}\big{[}p_{\theta}(x^{t:t+\delta}|z,x^{t-\delta:t})-\beta_{aux}(y^{\prime}-\hat{y}^{\prime})^{2}\big{]}$ (2)
$-\beta_{KL}KL\left[q_{\theta}(z|x^{t:t+\delta}){\mid\mid}p_{\theta}(z|y^{\prime})\right],$
@@ 3.1 Simulation Evaluation
Next to investigating the results qualitatively, we quantitatively measure the performance of VADA’s simulations and compare results to a state-of-the-art nanopore simulation approach that is not fully data-driven. We wish to compare simulated distribution to experimental nanopore current data, to quantify VADA’s effectiveness in capturing k-mer-specific variability. We utilize the Kolmogorov-Smirnov (KS) test statistic $D_{KS}$ as a metric to compute the similarity between the distribution of measurement values conditioned on the k-mer for simulated samples versus the corresponding distribution of experimental samples.
@@
3.1.1 Ablation Analysis
We further investigate the importance of the individual parts of architecture for good simulation performance by performing an ablation analysis. First, we evaluate an alternative to VADA that does not use a conditional prior. Instead, we use an unconditional gaussian $\mathcal{N}(0,1)$ and condition the decoder $p_{\theta}(x^{t:t+\delta}|z,x^{t-\delta:t},y^{\prime})$ and approximate posterior $q_{\theta}(z|x^{t:t+\delta},y^{\prime})$. Second, we evaluate VADA without the auxiliary regressor. Details on the training procedure can be found in Appendix 0.A. From the results in Table 1, we conclude that indeed conditioning via the prior is essential for achieving simulation performance competitive to the non-data-driven DeepSimulator. Furthermore, we conclude that including an auxiliary regressor does not clearly improve simulation performance and its use is only motivated by the requirement of learning an informative latent space.
Table 1: Evaluation of VADA and DeepSimulator in terms of simulation performance, measured by mean $\pm$ standard deviation of $D_{KS}$ (lower is better) over all k-mers.
Model Simulation Performance ($D_{KS}$)
DeepSimulator [11] $\mathbf{0.11}\pm 0.06$
VADA (unconditional prior) $0.32\pm 0.10$
VADA (no regressor) $\mathbf{0.13}\pm 0.04$
VADA (ours) $\mathbf{0.13}\pm 0.03$
@@ 3.2 Analysis of Latent Space
Furthermore, to investigate if the learned latent space can be used for the analysis of underlying sources of variation, we investigate if we can perform multi-label classification on the latent space. Specifically, we first sample a latent sample $z$ by using the approximate posterior, i.e. the encoder, where intuitively $z$ might contain a description of the sources of variation that resulted in nanopore current measurements $x^{t:t+\delta}$. Together with the embedding matrix $\theta_{f}$ of the embedding layer, $z$ is processed by a classifier $g(z,\theta_{f})$ to produce a binary class prediction for every possible k-mer. Specifically, there are $11025=4^{5}+$ 1 classes in total, representing all possible k-mer combinations of length $5$ and a class for incomplete k-mers, which can occur at the start or end of a nanopore current sequence. We measure performance by computing the Area Under the ROC (AUC) on the test set and report a resulting AUC of $\mathbf{0.91}$.We conclude that VADA has learned an informative latent representation that can be used for determining the k-mers that correspond to a window of nanopore measurements, thereby, successfully extracting one of the underlying sources of variation.
@@ 4 Conclusions
In this work, we proposed VADA, a fully data-driven approach for probabilistic simulation of nanopores. We evaluate VADA on experimental nanopore sequencing data and demonstrate VADA performs competitively with existing non-data-driven approaches. Furthermore, we show VADA has learned a meaningful latent representation by demonstrating we can accurately classify k-mers from this representation. To conclude, we demonstrate the value of deep generative models for simulating nanopore sequencing and capturing the underlying sources of variability for nanopore currents.

### v2

@@ Abstract
Nanopore sequencing offers the ability for real-time analysis of long DNA sequences at a low cost, enabling new applications such as early detection of cancer. Due to the complex nature of nanopore measurements and the high cost of obtaining ground truth datasets, there is a need for nanopore simulators. Existing simulators rely on handcrafted rules and parameters and do not learn an internal representation that would allow for analysing underlying biological factors of interest. Instead, we propose VADA, a purely data-driven method for simulating nanopores based on an autoregressive latent variable model. We embed subsequences of DNA and introduce a conditional prior to address the challenge of a collapsing conditioning. We experiment with an auxiliary regressor on the latent variable to encourage our model to learn an informative latent representation. We empirically demonstrate that our model achieves competitive simulation performance on experimental nanopore data. Moreover, we show our model learns an informative latent representation that is predictive of the DNA labels. We hypothesize that other biological factors of interest, beyond the DNA labels, can potentially be extracted from such a learned latent representation.
@@ 1 Introduction
To condition our model in accordance with the physical properties of nanopore sequencing, we represent the DNA sequence by embedding k-mers of DNA, as done in other machine learning tasks in this domain [16]. Initial empirical results showed that a straightforward approach to conditioning the latent variable model results in a conditioning collapse, where the model ignores the conditioning and produces bad samples unrelated to the DNA sequence. To overcome this problem, and effectively condition our model on a sequence of DNA, we introduce a conditional prior distribution on the latent space.
@@ 2.1.2 DNA representation
Multiple DNA bases influence each nanopore current measurement. Consequently, as mentioned earlier, we represent $y_{t}$ as a k-mer containing $k=5$ DNA bases. The embedding layer $f_{\theta}$ processes each $y_{t}$ in the window independently, producing embeddings $e^{t:t+\delta}=f_{\theta}(y^{t:t+\delta})$.
@@ 2.1.4 Informative latent space
Our model inherently learns a latent representation of the nanopore currents via the latent variable $z$ and is further encouraged to encode information about the DNA sequence into $z$ through the use of the conditional prior [7]. Nevertheless, there is no guarantee that $z$ captures the true underlying sources of variation, such as the DNA sequence that was in the pore.
To show that this is indeed the case, we experiment with a modified version of our model which is encouraged tolearn a representation that contains information about the DNA sequence through use of an auxiliary regressor during training, as done in other VAE’s that aim to learn an informative latent space, such as DIVA [7]. Specifically, during training, this regressor processes a latent sample $z$ to predict the aggregated embedding, $y^{\prime}$, that was used to condition the prior, as $\hat{y}^{\prime}=r_{\theta}(z)$. This auxiliary regressor is included in the visualization of the training procedure (Fig. 3).
In Section 3.2, we show that the model without this additional regressor performs as well as the model with the added regression objective, showing that the conditional prior indeed learns informative latent representations of the various k-mers.
Figure 3: VADA training overview, including the additional regressor used for experimentation in Section 3.2.
@@ 2.1.5 Training
Following the VAE framework [9] we utilize an approximate posterior $q_{\theta}(z|x^{t:t+\delta})$ during training. We optimize our model using a modified $\beta$-VAE loss [6, 7]. In our experiments with the auxiliary regressor, Mean Squared Error (MSE) is used to optimize the regressor. The contribution of the MSE to the overall loss is scaled using a hyperparameter $\beta_{\text{aux}}$. The complete loss term for VADA is given by Equation 2 and the loss in the experiments with the auxiliary regressor is given by Equation 3.
$\mathcal{L}_{\mathrm{VADA}}(x^{t:t+\delta},x^{t-\delta:t},y^{t:t+\delta})={}$ $-\mathbb{E}_{q_{\theta}(z\mid x^{t:t+\delta})}\big{[}\log p_{\theta}(x^{t:t+\delta}\mid z,x^{t-\delta:t})\big{]}$ (2)
$+\beta_{KL}\,D_{KL}\left[q_{\theta}(z\mid x^{t:t+\delta})\,{\mid\mid}\,p_{\theta}(z\mid y^{\prime})\right],$
$\mathcal{L}_{aux}(x^{t:t+\delta},x^{t-\delta:t},y^{t:t+\delta})={}$ $\mathcal{L}_{\mathrm{VADA}}(x^{t:t+\delta},x^{t-\delta:t},y^{t:t+\delta})$ (3)
$+\mathbb{E}_{q_{\theta}(z|x^{t:t+\delta})}\big{[}\beta_{aux}(y^{\prime}-\hat{y}^{\prime})^{2}\big{]}$
@@ 3.1 Simulation Evaluation
Besides investigating the results qualitatively, we quantitatively measure the performance of VADA’s simulations and compare the results to a state-of-the-art nanopore simulation approach that is not fully data-driven. We wish to compare simulated distributions to experimental nanopore current data, to quantify VADA’s effectiveness in capturing k-mer-specific variability. We use the Kolmogorov-Smirnov (KS) test statistic $D_{KS}$ as a metric to compute the similarity between the distribution of measurement values conditioned on the k-mer for simulated samples versus the corresponding distribution of experimental samples.
@@
Table 1: Evaluation of VADA and DeepSimulator in terms of simulation performance and base classification performance. We report mean $\pm$ standard deviation of $D_{KS}$ (lower is better) over all k-mers and AUC (higher is better). For each metric, we report score (standard deviation) over five independent runs.
Model Simulation Performance ($D_{KS}$) Base Classification (AUC)
DeepSimulator [11] $\mathbf{0.10}\ (0.000)\pm 0.06\ (0.000)$
VADA (unconditional prior) $0.32\ (0.001)\pm 0.10\ (0.003)$ $0.82\ (0.002)$
VADA (ours) $\mathbf{0.13}\ (0.003)\pm 0.03\ (0.002)$ $\mathbf{0.92}\ (0.000)$
VADA (with regressor) $\mathbf{0.13}\ (0.002)\pm 0.03\ (0.001)$ $\mathbf{0.92}\ (0.001)$
@@ 3.2 Analysis of Latent Space
Furthermore, to investigate if the learned latent space can be used for the analysis of underlying sources of variation, we investigate if we can perform multi-label classification on the latent space. Specifically, we first sample a latent sample $z$ by using the approximate posterior, i.e. the encoder, where intuitively $z$ might contain a description of the sources of variation that resulted in nanopore current measurements $x^{t:t+\delta}$. Together with the embedding matrix $\theta_{f}$ of the embedding layer, $z$ is processed by a classifier $g(z,\theta_{f})$ to produce a binary class prediction for every possible k-mer.
There are $11025=4^{5}+$ 1 classes in total, representing all possible k-mer combinations of length $5$ and a class for incomplete k-mers, which can occur at the start or end of a nanopore current sequence. We measure performance by computing the Area Under the ROC (AUC) on the test set and report a resulting AUC of $\mathbf{0.92}$.
We conclude that VADA has learned an informative latent representation that can be used for determining the k-mers that correspond to a window of nanopore measurements, thereby, successfully extracting one of the underlying sources of variation.
3.2.1 Importance of the conditional prior
To show that the conditional prior is indeed the right tool for creating an informative latent space, we perform an ablation analysis where we compare to a model that does not use a conditional prior, and to one that, on top of using a conditional prior, is trained together with an auxiliary regressor to potentially make the latent-space even more informative for basecalling.
In the model with the unconditional prior, we use an unconditional gaussian $\mathcal{N}(0,1)$ and condition the decoder $p_{\theta}(x^{t:t+\delta}|z,x^{t-\delta:t},y^{\prime})$ and approximate posterior $q_{\theta}(z|x^{t:t+\delta},y^{\prime})$ on the aggregated k-mer embedding $y^{\prime}$, instead.
To train the model with the auxiliary regressor, we use the loss shown in (3).
Details on the training procedure can be found in Appendix 0.A. From the results in Table 1, we conclude that indeed conditioning via the prior is essential for achieving simulation performance competitive to the non-data-driven DeepSimulator. Neglecting the prior in the conditioning mechanism clearly results in worse simulation performance and in a much less interpretable latent-space. On the otherhand, using an auxiliary regressor to further encourage an informative latent-space seems to provide no added value over using just the conditional prior.
@@ 4 Conclusions
In this work, we proposed VADA, a fully data-driven approach for probabilistic simulation of nanopores. We evaluate VADA on experimental nanopore sequencing data and demonstrate that VADA performs competitively with existing non-data-driven approaches. Furthermore, we show VADA has learned a meaningful latent representation by demonstrating that we can accurately classify k-mers from this representation. To conclude, we demonstrate the value of deep generative models for simulating nanopore sequencing and capturing the underlying sources of variability for nanopore currents.